OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write two equations which represent the two sentences.

OpenStudy (dumbcow):

x+x+y = 46 _> 2x+y =46 y = x + 4

OpenStudy (anonymous):

Ivy, check the other triangle question you asked, I have the answer there.

OpenStudy (anonymous):

:P